# No Detail Left Behind: Revisiting Self-Retrieval for Fine-Grained Image Captioning

### v1

@@ 1 Introduction
Image captioning, or generating natural language image descriptions, has witnessed remarkable progress over the last decade. Today’s captioning systems are composed of sophisticated deep learning architectures (Mokady et al., 2021, Stefanini et al., 2022, Dai et al., 2024, Liu et al., 2023) trained on vast datasets (Lin et al., 2014, Sharma et al., 2018, Thomee et al., 2016, Desai et al., 2021, Schuhmann et al., 2022). However, even with these advances, approaches often generate generic captions that are unable to differentiate similar images (see Figure 1), violating the fundamental purpose of a caption: to facilitate accurate and efficient communication of visual content (Fisch et al., 2020, Kreiss et al., 2022, Dessì et al., 2023). We attribute the shortcomings of image captioning systems to three key factors: (i) the nature of their training data, (ii) captioning evaluation metrics, and (iii) the maximum likelihood estimation (MLE) training approach.
@@
Training datasets may be divided into two: curated datasets (COCO (Lin et al., 2014), Flickr30k (Plummer et al., 2015)) or large-scale alt-text data (e.g. CC3M (Sharma et al., 2018)). While alt-text is noisy and may be unaligned with the image, human-annotated captions (COCO) may be generic (Kornblith et al., 2023) and lack world knowledge (Bavishi et al., 2023) as annotators describe visual concepts in a simplistic manner (e.g. labeling a Golden Retriever as a dog). Recently, foundation models are being used to enhance large-scale alt-text data by making them denser (Doveh et al., 2024, Urbanek et al., 2024). However, such methods are prone to inherit biases (e.g. gender, geography) present in foundation models (Hall et al., 2023, Sirotkin et al., 2022, Basu et al., 2023, Salman et al., ) and exhibit verbose language modeling priors (Liu et al., 2024).
To address these data inadequacies, we propose Visual Caption Boosting (VCB), a model agnostic framework designed to generate dense captions that holistically capture different aspects of the image (objects, attributes, relations, scene, etc.) while remaining anchored in human annotations (see Section 2). In brief, multiple human annotated captions are blended together using a Large Language Model (LLM) and expanded with an image description generated from a Multimodal Large Language Model (MLLM). To prevent conflicting visual details, we prompt the LLM to prefer the blended caption over the description from the MLLM. Thus, VCB creates fine-grained captions that are grounded in human annotations, enabling rich and informative datasets to train image captioning systems.
Figure 1: For a set of similar images, captioning systems struggle to generate meaningful captions that uniquely describe each image. COCO MLE: A model trained on COCO with MLE generates the same description for all images. COCO SR (Dessì et al., 2023): While the SR objective may help, the COCO captions are not rich enough and lead to hallucinations such as “two people” (middle). OURS SR: Our improved visual captions and SR fine-tuning with hard negatives results in discriminative captions.
@@ II. Image caption evaluation
metrics can be broadly classified into reference-based and reference-free. Reference-based metrics such as BLEU (Papineni et al., 2002), CIDEr (Vedantam et al., 2015), and SPICE (Anderson et al., 2016) rely on comparisons to ground-truth (GT) captions that may not include image details beyond salient objects. These metrics tend to penalize captions that are more specific than the ground-truth, often favoring generic descriptions (Wang et al., 2020). Reference-free metrics like CLIPScore (Hessel et al., 2021) alleviate this issue by directly measuring image-text similarity, but may fail to encourage diverse and discriminant captions. These limitations necessitate the development of new evaluation strategies that incentivize models to produce holistic and distinct captions.
In this work, we ask what makes a “good” description when evaluating captioning systems? We posit that a caption should enable a listener to pick out a target image from a bag of images with similar visual elements (Liu et al., 2018). To this extent, we propose to evaluate captioning systems through the lens of self-retrieval (SR) i.e., their ability to retrieve an image using its generated caption against of bag of highly similar distractor images. Traditionally, improvements in text-to-image (T2I) retrieval have focused on either enhancing the embedding spaces of scoring functions (Faghri et al., 2017) or recently, making denser, more detailed text queries (Lai et al., 2023, Li et al., 2024, Singla et al., 2024, Shvetsova et al., 2023). In this work, we introduce a novel perspective: by curating image sets of highly similar images, we can use T2I retrieval to assess whether the query (caption) captures fine-grained visual concepts necessary to retrieve the correct image. Specifically, we construct bags of highly similar images, where uniquely describing each image within a bag requires describing various facets of visual understanding. For example, retrieving the target image from Figure 1 requires the caption to incorporate information about attributes (red bike) or orientation (inverted body). Hence, we evaluate whether the captioning system captures subtle visual distinctions by retrieving the target image using its generated caption without looking at the other candidate images in the bag.
@@
MLE training incentivizes captioning systems to overuse common concepts and statistically probable phrases when describing visually similar images. While previous works have optimized for self-retrieval (SR) reward to learn discriminant captioners (Luo et al., 2018, Liu et al., 2018, Rennie et al., 2017, Dessì et al., 2023), the models are first trained via MLE on generic caption datasets (e.g. COCO). This is suboptimal and we show that SR fine-tuning is sensitive to initialization: it is necessary to start with a captioning system that captures fine-grained details in order to better preserve faithfulness to GT captions. In fact, we discover a propensity of captioning systems to hallucinate (Section 4.4) when trained via SR on generic captions (Dessì et al., 2023). Thus, current SR approaches face two distinct challenges: a trade-off between retrieval performance and caption faithfulness, and sub-optimal fine-grained retrieval performance.
To enhance SR’s ability to instill fine-grained visual information, we (i) fine-tune both components (language and visual) of the captioning system (Section 4.5), and (ii) mine multiple hard negatives (visually similar sets of images) to create training bags. We also introduce a curriculum learning approach that progressively increases the bag size during training to leverage the contrastive nature of our retrieval-based reward (Section 4.6). Our carefully designed curriculum enables SR to leverage the rich initialization provided by VCB, improving retrieval performance as well as the faithfulness of the generated captions. Through our training procedure, we are able to circumvent the aforementioned trade-off and achieve substantial performance gains, surpassing (Dessì et al., 2023) by 20% on TrueMatch without making any modifications to the model architecture or reward. Our results demonstrate that a well-crafted training paradigm achieves performance comparable to MLLMs that are orders of magnitude larger. Figure 1 shows images of a bag from TrueMatch and captions generated by the same model architecture trained with different datasets (COCO vs. Ours) and paradigms (MLE vs. vanilla SR vs. Ours SR).
@@
We address three challenges plaguing current image captioning systems: Data, Evaluation, and the MLE training. (i) We begin by identifying the inadequacies of image captioning datasets in Section 2, and propose Visual Caption Boosting, a novel caption enhancement strategy that leverages LLMs and MLLMs to generate dense, informative, and unbiased captions that are anchored in human annotations. (ii) We then introduce TrueMatch (Section 3), a benchmark consisting of curated bags of highly similar images that uses self-retrieval (SR) to assess the ability of captioning systems to capture fine-grained visual distinctions. (iii) Section 4 leverages the rich data and effective evaluation methods introduced in the previous two sections to train fine-grained captioning systems with SR. We offer new insights into SR fine-tuning through extensive ablations revealing its sensitivity to MLE initialization, failure to preserve caption faithfulness, and tendency to hallucinate. (iv) Finally, we design a simple “plug and play” training recipe that enables SR to improve caption faithfulness while significantly outperforming vanilla SR (Dessì et al., 2023) on TrueMatch and achieving state-of-the-art results on ImageCoDe (Krojer et al., 2022), another benchmark for fine-grained image retrieval.
@@ 2 Improving Datasets through Visual Caption Boosting
Figure 2: Example of Visual Caption Boosting: transforming the original human annotated captions to HolisticCap. First, an LLM blends the human annotations to create BlendCap. Next, an MLLM generates a dense visual caption that may be noisy. Finally, we create HolisticCap by instructing the LLM to incorporate fine-grained details from the Visual Caption with the BlendCap, while staying anchored in human annotations in case of conflicts. Specific prompts are shared in Section A.1. The colors indicate various concepts extracted from the human annotations or the visual caption. The red underlined text, indicating hallucinations or verbose text, is ignored in the holistic caption as we anchor the visual caption to human annotations.
@@
To address this, recent works leverage foundation models to synthetically expand visual information within captions (Doveh et al., 2024, Li et al., 2024, Singla et al., 2024, Lai et al., 2023), instilling a wider range of visual details. However these methods are prone to inherit biases present in foundation models (Sirotkin et al., 2022, Basu et al., 2023, Salman et al., ). Moreover, the descriptions generated by some of these methods are excessively long making them susceptible to hallucinations (Favero et al., 2024). They also pose challenges to training captioning systems as they exceed the token capacity of current VLMs (e.g. CLIP with 77 tokens) (Urbanek et al., 2024). Nevertheless, while the individual COCO captions are sparse, we find they describe complementary facets of the image, e.g. “watering hole”, “bird flying over”, “herd of zebras” (Figure 2) that may provide additional details. This also corroborates findings by Ye et al. (2023) that annotators from different cultural backgrounds describe different visual concepts when viewing the same image.
@@
Building upon these findings, we introduce Visual Caption Boosting (VCB), a novel two-stage approach to enrich the training data with dense, more informative captions that encourage captioning systems to learn and generate rich descriptions. VCB leverages foundation models and the diverse perspectives offered by human annotators to generate rich descriptions while being anchored in human data.
BlendCap
@@
leverages an off-the-shelf LLM to create a blended caption that combines multiple facets of visual information that the human annotators describe. Figure 2 shows an example of how a few captions can be blended together into a comprehensive description of the image using our method. Notably, we prompt the LLM to minimize redundant information resulting in short descriptions.
HolisticCap
@@
builds upon BlendCap by incorporating a fine-grained visual description produced by an MLLM. Specifically, we prompt the LLM to instill the visual caption into the blended caption while preferring human-grounded BlendCap in case of conflicting visual information. As seen in Figure 2, this enables the LLM to ignore specific details from the visual caption such as “river or lake” and “drinking”, as they conflict with BlendCap’s description like “watering hole” and “grazing and standing”. Additionally, the anchoring of semantic visual information in human annotations encourages the LLM to eliminate verbose tendencies of MLLMs, producing rich and succinct captions that capture fine-grained details. Although recent works (Urbanek et al., 2024, Lai et al., 2023) address similar challenges, they have some drawbacks as they anchor synthetic visual information in alt-text (Lai et al., 2023) or depend on manual annotation for incorporating dense visual information (Urbanek et al., 2024).
@@
In summary, Visual Caption Boosting is able to efficiently instill fine-grained visual information into image caption datasets while being anchored in human annotations. Section A.1 presents the LLM and MLLM prompts and Section A.2 analyzes the caption lengths distributions.
@@ 3 TrueMatch: Fine-grained evaluation through Self-Retrieval
Within a bag of images $\mathcal{B}$ (or a minibatch during training), we require the generated caption $c$ for an image $i$ to retrieve itself (image $i$) from $\mathcal{B}$. Note, $\mathcal{B}$ contains $i$ and a set of visually similar distractor images $\mathcal{D}$. The caption $c$ and all images in $\mathcal{B}$ are encoded using CLIP text and vision encoders and ranking is performed by computing the cosine similarity $\text{sim}(c,i)$ in the CLIP embedding space. Consequently, the caption is deemed to be good $\iff\text{sim}(c,i)>\text{sim}(c,i^{\prime}),\,\forall i^{\prime}\in D$.
@@ 3.1 Benchmark Creation
Given a dataset, the benchmark creation process involves curating bags of highly similar images. We use 10,000 images from COCO’s validation and test sets (Karpathy & Fei-Fei, 2015) for our benchmark.
@@
process is performed to retain only those bags that would require captioning systems to capture some aspect of fine-grained visual discrimination. For example, for bag size 3, this reduces the number of bags in TrueMatch from 680 to 254.
@@ 3.2 Experiment 1: Results on TrueMatch
We evaluate several open-source captioning approaches, MLLMs, and SR trained models on TrueMatch. Recall@1 is reported in Table 1 for bag sizes 3, 5, and 7.
@@
Irrespective of their size, captioning approaches struggle to capture fine-grained visual details leading to poor performance on TrueMatch.

### v2

@@ 1 Introduction
Image captioning, or generating natural language image descriptions, has witnessed remarkable progress over the last decade. Today’s captioning systems are composed of sophisticated deep learning architectures (Mokady et al., 2021, Stefanini et al., 2022, Dai et al., 2024, Liu et al., 2023) trained on vast datasets (Lin et al., 2014, Sharma et al., 2018, Thomee et al., 2016, Desai et al., 2021, Schuhmann et al., 2022). However, even with these advances, approaches often generate generic captions that are unable to differentiate similar images (see Figure 1), violating the fundamental purpose of a caption: to facilitate accurate and efficient communication of visual content (Fisch et al., 2020, Kreiss et al., 2022, Dessì et al., 2023, Das et al., 2017). We attribute the shortcomings of image captioning systems to three key factors: (1) the nature of their training data, (2) captioning evaluation metrics, and (3) the maximum likelihood estimation (MLE) training approach.
@@
Training datasets may be divided into two: curated datasets (COCO (Lin et al., 2014), Flickr30k (Plummer et al., 2015)) or large-scale alt-text data (e.g. CC3M (Sharma et al., 2018)). While alt-text is noisy and may be unaligned with the image, human-annotated captions (COCO) may be generic (Kornblith et al., 2023) and lack world knowledge (Bavishi et al., 2023) as annotators describe visual concepts in a simplistic manner (e.g. labeling a Golden Retriever as a dog). Recently, foundation models are being used to enhance large-scale alt-text data by making them denser (Doveh et al., 2024, Urbanek et al., 2024). However, such methods are prone to inherit biases (e.g. gender, geography) present in foundation models (Hall et al., 2023, Sirotkin et al., 2022, Basu et al., 2023, Salman et al., 2022) and exhibit verbose language modeling priors (Liu et al., 2024).
To address these data inadequacies, we propose Visual Caption Boosting (VCB), a model agnostic framework designed to generate dense captions that holistically capture different aspects of the image (objects, attributes, relations, scene, etc.) while remaining anchored in human annotations (see Section 2). In brief, multiple human annotated captions are blended together using a Large Language Model (LLM) and expanded with an image description generated from a Multimodal Large Language Model (MLLM). In case of conflicting visual details, we prompt the LLM to prefer the blended caption over the description from the MLLM. Thus, VCB creates fine-grained captions that are grounded in human annotations, enabling rich and informative datasets to train image captioning systems.
Figure 1: For similar images, captioning systems struggle to generate meaningful captions that uniquely describe each image. In this example, COCO MLE: A model trained on COCO with MLE generates the same generic caption for the first two images. COCO SR (Dessì et al., 2023): While the self-retrieval (SR) objective may help, the COCO captions are not rich enough to generate salient visual details. OUR SR: Our improved data and training recipe results in fine-grained, and therefore discriminant captions.
@@ II. Image Caption Evaluation
. Current metrics can be broadly classified into reference-based and reference-free. Reference-based metrics such as BLEU (Papineni et al., 2002), CIDEr (Vedantam et al., 2015), and SPICE (Anderson et al., 2016) rely on comparisons to ground-truth (GT) captions that may not include image details beyond salient objects. Although these metrics evaluate grammatical correctness, they tend to penalize captions that are more specific than the ground-truth, often favoring generic descriptions (Wang et al., 2020). Reference-free metrics like CLIPScore (Hessel et al., 2021) alleviate this issue by directly measuring image-text similarity, but may fail to assess discriminativeness of the captions. These limitations necessitate the development of new evaluation strategies that incentivize the models to produce fine-grained captions.
In this work, we ask what makes a “fine-grained” or “good” description when evaluating captioning systems? We posit that a caption should be succinct, but enable a listener to identify the target image from a bag of images with similar visual elements (Liu et al., 2018). To this extent, we propose to evaluate captioning systems through the lens of self-retrieval (SR). We measure the ability to retrieve the target image based on the generated caption within a bag of highly similar distractor images.
Traditionally, improvements in text-to-image (T2I) retrieval have focused on either enhancing the scoring functions (Faghri et al., 2017) or making denser, more detailed text queries (Lai et al., 2023, Li et al., 2024, Singla et al., 2024, Shvetsova et al., 2024). In this work, we introduce a novel perspective: by curating sets of similar images, we can use T2I retrieval as a tool for evaluating whether the text query (caption) captures fine-grained visual details necessary for retrieving the target image. Specifically, we construct bags of highly similar images, designed such that uniquely describing each image within a bag demands capturing different facets of visual understanding. For example, retrieving the middle image in Figure 1 requires the caption to incorporate information about attributes (blue frisbee) or background objects (fence). Note, each caption is generated without looking at the other images in the bag. Thus, we use discriminability as a proxy to measure how well a model describes fine-grained components in an image.
@@
MLE training incentivizes captioning systems to overuse common concepts and statistically probable phrases when describing visually similar images. While previous works have optimized self-retrieval (SR) to learn discriminant captioners (Luo et al., 2018, Liu et al., 2018, Cho et al., 2022, Dessì et al., 2023), the models are first trained via MLE on generic caption datasets (e.g. COCO). We find that this is suboptimal and show that SR fine-tuning is sensitive to initialization: it is necessary to start with a captioning system that captures fine-grained details in order to better preserve faithfulness to GT captions. In fact, we discover a propensity of captioning systems to hallucinate (Section 4.4) when trained via SR on generic captions (Dessì et al., 2023). Thus, current SR approaches face two distinct challenges: a trade-off between retrieval performance and caption faithfulness, and sub-optimal fine-grained retrieval performance.
To enhance SR’s ability to instill fine-grained visual information, we (1) fine-tune both components (language and visual) of the captioning system (Section 4.5), and (2) mine multiple hard negatives (visually similar sets of images) to create training bags. We also introduce a curriculum learning approach that progressively increases the bag size during training to leverage the contrastive nature of our retrieval-based reward (Section 4.6). Our carefully designed curriculum enables SR to leverage the rich initialization provided by VCB, improving retrieval performance as well as the faithfulness of the generated captions. Through our training procedure, we are able to circumvent the aforementioned trade-off and achieve substantial performance gains, surpassing (Dessì et al., 2023) by 20% on TrueMatch without making any modifications to the model architecture or reward. Our results demonstrate that a well-crafted training paradigm achieves performance comparable to MLLMs that are orders of magnitude larger. Figure 1 shows images of a bag from TrueMatch and captions generated by the same model architecture trained with different datasets (COCO vs. Ours) and paradigms (MLE vs. Vanilla SR vs. Our SR).
@@
We address three challenges plaguing current image captioning systems: data, evaluation, and MLE training. (1) We identify the inadequacies of image captioning datasets in Section 2, and propose Visual Caption Boosting, a novel caption enhancement strategy that leverages LLMs and MLLMs to generate dense, informative, and unbiased captions that are anchored in human annotations. (2) We introduce TrueMatch (Section 3), a benchmark consisting of curated bags of highly similar images that uses self-retrieval (SR) to assess the ability of captioning systems to capture fine-grained visual distinctions. (3) In Section 4, we leverage the rich data and effective evaluation introduced in the previous two sections to train fine-grained captioning systems with SR. We offer new insights into SR fine-tuning through extensive ablations revealing its sensitivity to MLE initialization, failure to preserve caption faithfulness, and tendency to hallucinate. (4) Finally, we design a simple “plug and play” training recipe that enables SR to improve caption faithfulness while significantly outperforming vanilla SR (Dessì et al., 2023) on TrueMatch and achieving state-of-the-art results on ImageCoDe (Krojer et al., 2022), another benchmark for fine-grained image retrieval.
@@ 2 Improving Datasets through Visual Caption Boosting
Figure 2: Example of Visual Caption Boosting transforming the original human annotated captions (left) to a Holistic Caption. First, an LLM blends the human annotations to create a Blended Caption. Next, an MLLM generates a dense visual caption that may be noisy. Finally, we create a Holistic Caption by instructing the LLM to incorporate fine-grained details from the Visual Caption with the Blended Caption, while staying anchored in human annotations in case of conflicts. Specific prompts are shared in Section A.1. The colors indicate various concepts extracted from the human annotations or the visual caption. The red underlined text (illustrated by us for ease of understanding) indicating hallucinations or verbose text, is ignored in the Holistic Caption as we anchor it to human annotations.
@@
To address this, recent works leverage foundation models to synthetically expand visual information within captions (Doveh et al., 2024, Li et al., 2024, Singla et al., 2024, Lai et al., 2023, Jiao et al., 2024), instilling a wider range of visual details. However these methods are prone to inherit biases present in foundation models (Sirotkin et al., 2022, Basu et al., 2023, Salman et al., 2022). Moreover, the descriptions generated by some of these methods are excessively long making them susceptible to hallucinations (Favero et al., 2024). They also pose challenges to training captioning systems as they exceed the token capacity of current VLMs (e.g. CLIP with 77 tokens (Urbanek et al., 2024)). Nevertheless, while individual COCO captions are sparse, we find that they describe complementary facets of the image e.g. “watering hole”, “bird flying over”, “herd of zebras” (Figure 2). This also corroborates findings of Ye et al. (2023) that annotators of different cultural backgrounds describe distinct visual concepts when viewing the same image.
@@
Building upon these findings, we introduce Visual Caption Boosting (VCB), a novel two-stage approach to enrich the training data with dense, more informative captions that encourage captioning systems to learn and generate rich descriptions. VCB leverages foundation models and the diverse perspectives offered by human annotators to generate rich descriptions while being anchored in human annotations.
Step 1. BlendCap
@@
leverages an off-the-shelf LLM to create a blended caption that combines multiple facets of visual information that the human annotators describe. Figure 2 shows an example of how a few captions can be blended together into a comprehensive description of the image using our method. Notably, we prompt the LLM to minimize redundant information resulting in succinct descriptions.
Step 2. HolisticCap
@@
extends BlendCap by incorporating a fine-grained visual description produced by an MLLM [^1] We adopted InstructBLIP for this work, see https://huggingface.co/Salesforce/instructblip-vicuna-7b. Specifically, we prompt the LLM to instill the visual caption into the blended caption while preferring human-grounded BlendCap in case of conflicting visual information. As seen in Figure 2, this enables the LLM to ignore hallucinations present in Visual Caption such as “river or lake” and “drinking”, as they conflict with BlendCap’s description like “watering hole” and “grazing and standing”. Additionally, the anchoring of semantic visual information in human annotations encourages the LLM to eliminate verbose tendencies of MLLMs, producing rich and succinct captions that capture fine-grained details. Although recent works (Urbanek et al., 2024, Lai et al., 2023) address similar challenges, they have some drawbacks as they anchor synthetic visual information in alt-text (Lai et al., 2023) or depend on manual annotation for incorporating dense visual information (Urbanek et al., 2024).
Anchored captions hallucinate less
. We conduct a human study to measure the degree of hallucinations in BlendCap and HolisticCap (see Section A.2). We find that BlendCap introduces negligible hallucinations and that anchoring VisualCap in human annotations significantly (41.6%) reduces hallucinations in HolisticCap.
@@
In summary, Visual Caption Boosting is able to efficiently instill fine-grained visual information into image caption datasets while being less prone to hallucinations. Section A.1 presents the LLM and MLLM prompts and Section A.3 analyzes the distribution of the lengths of the generated captions.
@@ 3 TrueMatch: Fine-grained Evaluation through Self-Retrieval
Within a bag of images $\mathcal{B}$ (or a minibatch during training), we require the generated caption $c$ for an image $i$ to retrieve itself (image $i$) from $\mathcal{B}$. Note, $\mathcal{B}$ contains $i$ and a set of visually similar distractor images $\mathcal{D}$. The caption $c$ and all images in $\mathcal{B}$ are encoded using CLIP text and vision encoders respectively and ranking is performed by computing the cosine similarity $\text{sim}(c,i)$ in the CLIP embedding space. Consequently, the caption is deemed to be good $\iff\text{sim}(c,i)>\text{sim}(c,i^{\prime}),\,\forall i^{\prime}\in\mathcal{D}$.
@@ 3.1 Benchmark Creation
Given a dataset, the benchmark creation process involves curating bags of highly similar images. We use the 10,000 images from COCO’s validation and test sets (Karpathy & Fei-Fei, 2015) for our benchmark.
@@
pass is performed to retain only those bags that would require captioning systems to capture some aspect of fine-grained visual discrimination. For bag size 3, this reduces the number of bags in TrueMatch from 680 to 254. Some examples of selected and discarded bags are shown in Figure 10.
@@ 3.2 Experiment 1: Results on TrueMatch
We evaluate several open-source captioning approaches, MLLMs, and SR trained models on TrueMatch. Recall@1 (R@1) is reported in Table 1 for bag sizes 3, 5, and 7.
@@
Captioning models, irrespective of their size, struggle to capture fine-grained visual details leading to poor performance on TrueMatch.